OpenStudy (anonymous):

can this be symplify? please help x=2/(2√(2x+1))

OpenStudy (radar):

\[1\over \sqrt{2x+1}\]

OpenStudy (cruffo):

How about rationalizing the denominator?

OpenStudy (cruffo):

oh.... hold on ... didn't see the x on the left hand side of the equal sign.

OpenStudy (radar):

\[\sqrt{2x+1}\over 2x+1\] I did try that but it don't look any simpler lol

OpenStudy (anonymous):

Squaring you get, $$x^2(2x+1)=1$$ or $$2x^3+x^2-1=0$$

OpenStudy (cruffo):

There should be at least one real root between 0 and 1 :)

OpenStudy (radar):

You're right it don't check out

OpenStudy (radar):

Is it to be solved?....or to be simplified? Or both!

OpenStudy (cruffo):

If it is to be solved, you'll need Cardano's formula for cubics, or a numerical method like Bisection or Newton's. Wolfram Alpha shows a numerical root at around 0.6, and two imaginary roots.

OpenStudy (radar):

It is above my paygrade, I need a nap.

OpenStudy (cruffo):

lol

OpenStudy (radar):

I looked at the link, interesting one real root, and 2 imaginary.

OpenStudy (anonymous):

Thank you all that was very helpful